OpenStudy (anonymous):

Find the component form and magnitude of the vector 3AB A=(2,-1) B=(3,1)

OpenStudy (gorv):

AB=((3-2),(1-(-1)))=(1,1+1)=(1,2) magnitude AB=\[\sqrt{1^2+2^2}=\sqrt{5}\]

OpenStudy (gorv):

3AB =3 root 5